# Online Multi-Task Offloading for Semantic-Aware Edge Computing Systems

### v1

@@ IV-B QoE Performance and Convergence
3) QoE Preference: We can adjust the model to achieve desired results by modifying QoE preferences for different metrics. Fig. 10 shows the performance of latency and task accuracy under different weights by adjusting the weights assigned to latency and accuracy. As the weight for latency increases, tasks are offloaded to the ES with more powerful processing capabilities, and there tends to be a preference for a smaller semantic extraction factor, $\mu_{n}$, to maximize the compression rate. This results in a decrease in the accuracy of tasks executed by the ES. Fig. 11 presents the performance of energy consumption and task accuracy under different weights. A significant jump occurs at (0.67, 0.33) when the weight for accuracy increases, as tasks are offloaded to the ES, causing a substantial increase in energy consumption and a corresponding rise in accuracy. Fig. 12 describes the performance of energy consumption and latency under varying weights. When the weight for latency increases beyond (0.3, 0.7), latency begins to decrease, and energy consumption increases. This is due to an increase in the number of tasks offloaded to the ES, and users tend to choose a smaller semantic extraction factor, $\mu_{n}$, to reduce data transmission volume. However, a smaller $\mu_{n}$ means that users need more computational resources to extract more compact semantic information.
TABLE V: Performance under different preferences
Delay (ms) Energy Consumption (J) Accuracy (%)
No Preference 2.82 3.19 64.37
Delay Preference 2.18 2.66 54.08
Energy Preference 4.95 0.83 55.66
Accuracy Preference 3.25 3.49 68.57
Table V shows the performance in execution latency, energy consumption, and task accuracy under varying preference settings. Specifically, we adjust the relative magnitude of $\omega_{t}$, $\omega_{e}$, and $\omega_{a}$ to set preferences for different metrics and retrain to obtain the corresponding model for each. 1) For the delay-preference model, we set $\omega_{t}=1$, $\omega_{e}=0$ and $\omega_{a}=0$. The delay-preference model tends to choose a smaller $\mu_{n}$ to reduce the amount of data transmitted. This results in a partial loss of semantic information, causing a decline in task accuracy. 2) For the energy-preference model, we set $\omega_{t}=0$, $\omega_{e}=1$ and $\omega_{a}=0$. The energy-preference model favors local execution, leading to significantly higher execution latency while reducing energy consumption. However, task accuracy is unsatisfactory. 3) For the accuracy-preference model, we set $\omega_{t}=0$, $\omega_{e}=0$ and $\omega_{a}=1$. The accuracy-preference model aims to maximize the accuracy of all tasks, prompting it to choose a larger $\mu_{n}$ to preserve semantic information better. Nevertheless, transmitting a high volume of data results in higher overall energy consumption.
@@ V Conclusion
In this paper, we proposed a new multi-modal task offloading system and designed a unified quality of experience (QoE) metric to measure overall task offloading performance. Text classification, image classification, and visual question answering (VQA) tasks can be executed locally or offloaded to the edge server after semantic extraction. A multi-agent PPO reinforcement learning algorithm is used to solve the problem of maximizing overall user QoE through the joint optimization of the semantic extraction factor, computing resources, and communication resources. Experimental results show that the proposed method outperforms other benchmarks in overall user QoE. Specifically, the proposed algorithm achieves a reduction in execution latency and energy consumption of 18.1% and 12.9%, respectively compared to the semantic-unaware method. Additionally, the proposed approach can be flexibly adapted to user preferences, fully satisfying latency, energy, and task performance preference requirements.

### v2

@@ V Conclusion
In this paper, we proposed a new multi-task offloading system and designed a unified quality of experience (QoE) metric to measure overall task offloading performance. Text classification, image classification, and visual question answering (VQA) tasks can be executed locally or offloaded to the edge server after semantic extraction. A semantic-aware multi-agent PPO reinforcement learning algorithm is used to solve the problem of maximizing overall user QoE through the joint optimization of the semantic extraction factor, computing resources, and communication resources. Experimental results show that the proposed method outperforms other benchmarks in overall user QoE. Specifically, the proposed algorithm demonstrates a 12.68% improvement in QoE compared to the semantic-unaware approach. Additionally, the proposed approach can be flexibly adapted to user preferences, fully satisfying latency, energy, and task performance preference requirements.